1.Given the following information about a closed economy, what is the level of investment spending and national saving, and what is the budget balance? There are no government transfers.

GDP: $1,000 million T=$50 million

C=$850 million G=$100 million

2.Given the following information about an open economy, what is the level of investment spending and National savings, and what are the budget balance and net capital inflow? There are no government transfers.

GDP: $1,000 million G=$100 million

C=$850 million X=$100 million

T=$50 million IM=$125 million

3. If the multiplier equals 10, then what is the marginal propensity propensity to consume?
Answer #1

Answer:

1]

GDP= $1000 million

T= $50 million

C= $850 million

G = $100 million

I = Investment= GDP- C-G

I = 1000-850-100

I= $50 milion

Reason - when we calculate GDP by expenditure method GDP is equal to the sum of consumption expenditure, Investment expenditure, and Government spending.

Private Savings= GDP - T- C

Private Savings= 1000 -50 -850

Private Savings= $100 million

Reason- Private savings are the savings of the household. They are calculated by subtracting Taxes and consumption expenditure from the income or GDP.

Government savings or Public savings = T - G

Government savings or Public savings = 50- 100 = -$50 million

Reason- Public Savings are the savings of the government. they are calculated by subtracting the expenditure from the taxes.

Since Government savings is positive its Budget surplus of $50 million.

National Savings= Private savings + Public Savings

National Savings= $100 million - $50 million = $50 million

Reason- National savings is the sum of savings with the government and savings with the private households.

As per HomeworkLib policy 1 question per post.